# Impact of Dataset Properties on Membership Inference Vulnerability of Deep Transfer Learning

### v5

@@ Abstract
Membership inference attacks (MIAs) are used to test practical privacy of machine learning models. MIAs complement formal guarantees from differential privacy (DP) under a more realistic adversary model. We analyse MIA vulnerability of fine-tuned neural networks both empirically and theoretically, the latter using a simplified model of fine-tuning. We show that the vulnerability of non-DP models when measured as the attacker advantage at a fixed false positive rate reduces according to a simple power law as the number of examples per class increases. A similar power-law applies even for the most vulnerable points, but the dataset size needed for adequate protection of the most vulnerable points is very large.
@@ 1 Introduction
To achieve our goal, we theoretically analyze and systematically apply two state-of-the-art black-box MIAs, LiRA (Carlini et al., 2022) and RMIA (Zarifzadeh et al., 2024), to help understand practical privacy risks when fine-tuning deep-learning-based classifiers without DP protections. For the theoretical model that we analyse, LiRA is the optimal attack by the Neyman–Pearson lemma (Neyman and Pearson, 1933). Under the black-box threat model, in which the adversary does not have access to the model parameters, LiRA and RMIA have been shown to empirically outperform other attacks, especially when the number of shadow models is sufficiently large.
@@
Related work There has been evidence that classification models with more classes are more vulnerable to MIA (Shokri et al., 2017), models trained on fewer samples can be more vulnerable (Chen et al., 2020; Németh et al., 2025), and classes with less examples tend to be more vulnerable (Chang and Shokri, 2021; Kulynych et al., 2022; Tonni et al., 2020). A larger generalisation error, which is related to dataset size, has also been shown to be sufficient for MIA success (Song and Mittal, 2021), though not necessary (Yeom et al., 2018). Similarly, minority subgroups tend to be more affected by DP (Suriyakumar et al., 2021; Bagdasaryan et al., 2019). Feldman and Zhang (2020) showed that neural networks trained from scratch are required to memorize a significant fraction of their training data to obtain high utility, while the memorisation is greatly reduced for fine-tuning. Additionally, Tobaben et al. (2023) reported how the MIA vulnerability of few-shot image classification is affected by the number of shots. Yu et al. (2023) studied the relationship between the MIA vulnerability and individual privacy parameters for different classes. Recently, worst-case MIA vulnerability has gained more attention (Guépin et al., 2024; Meeus et al., 2024; Azize and Basu, 2024). Nonetheless, the prior works do not consider the rate of change in the vulnerability evaluated at a low fpr, as dataset properties change. Our work significantly expands on these works by a) explicitly identifying a quantitative relationship between dataset properties and MIA vulnerability, i.e., the power-law in Equation˜1, and b) focusing on the worst-case vulnerability, both evaluated at a low FPR. This in turn allows us to extrapolate MIA vulnerability to DP guarantee.
@@
    MIA experiments on the average case vulnerability: We conduct a comprehensive study of MIA vulnerability (tpr at a fixed low fpr) in the transfer learning setting for image classification tasks with target models trained using many different datasets with varying sizes and confirm the theoretical power-law between the number of examples per class and the vulnerability to MIA (see Figures˜1 and 4.2). We fit a regression model which follows the functional form of the theoretically derived power-law. We show both a very good fit on the training data as well as a good prediction quality on unseen data from a different feature extractor and when fine-tuning other parameterisations (see Section˜4.3).
@@ 3.4 A simplified model of the optimal membership inference
We sample vectors on a high-dimensional unit sphere and classify them based on inner product with estimated class mean. This model is easier to analyse theoretically than real-world deep learning examples. We generate the data and form the classifiers (which are our target models) as follows:
@@
    For each class, we first sample a true class mean $m_{c}$ on a high dimensional unit sphere that is orthogonal to all other true class means ($\forall i,j\in\{1,\dots,C\},i\neq j:m_{i}\perp m_{j}$).
@@
    We sample $2S$ vectors $\bm{x}_{c}$ for each class. We assume that they are Gaussian distributed around the the true class mean $\bm{x}_{c}\sim\mathcal{N}(m_{c},\Sigma)$ where the $\Sigma$ is the in-class covariance.
@@
    For each ”target model” we randomly choose a subset of size $CS$ from all generated vectors and compute per-class means $\hat{m}_{c}$.
@@
    The computed mean is used to classify sample $\bm{x}$ by computing the inner product $\langle\bm{x},\hat{m}_{c}\rangle$ as a metric of similarity.
@@
$\log(\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))\\ =-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}+\log(1+\xi(S)),$ (9)
$\log(\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))\approx-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}.$ (10)
@@
This implies that if $\bm{x}-m_{\bm{x}}||$ is bounded, then the worst-case vulnerability is also bounded. Hence we can significantly reduce the MIA vulnerability of all examples in this non-DP setting by simply increasing the number of examples per class.
@@
By Lemma˜1 and the proof of Theorem˜2, a necessary condition for the power-law is that $(\mu_{\mathrm{in}}(\bf{x})-\mu_{\mathrm{out}}(\bf{x}))/\sigma(\bf{x})$ converges to zero at rate $O(1/S^{\alpha})$ with $\alpha>0$. In our simplified model, this holds with $\alpha=1/2$. However, $\mu_{\mathrm{in}}(\bf{x})-\mu_{\mathrm{out}}(\bf{x})\rightarrow 0$ would not always be the case for larger neural networks trained from scratch. Therefore, we do not expect similar power-laws in more general training algorithms.
@@
$\log(\overline{\textsc{tpr}}_{\mathrm{LiRA}}-\overline{\textsc{fpr}}_{\mathrm{LiRA}})\approx-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\overline{\textsc{fpr}}_{\mathrm{LiRA}})^{2}+\log\left(\mathbb{E}_{(\bm{x},y)\sim{\mathbb{D}}}\left[\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}\right]\right).$ (12)
@@ 4.1 Experimental setup
We report results for fine-tuning a last layer classifier (Head) trained on top of a Vision Transformer ViT-Base-16 (ViT-B; Dosovitskiy et al., 2021), pre-trained on ImageNet-21k (Russakovsky et al., 2015). The results for using ResNet-50 (R-50; Kolesnikov et al., 2020) as a backbone can be found in Section˜D.1. We optimise the hyperparameters (batch size, learning rate and number of epochs) using the Optuna library (Akiba et al., 2019) with the Tree-structured Parzen Estimator (TPE; Bergstra et al., 2011) sampler with 20 iterations (more details in Section˜C.2). We provide the the code for reproducing the experiments in an open repository [^1] https://github.com/DPBayes/impact-dataset-properties-MI-vulnerability-deep-TL.
@@ 4.3 Model to predict dataset vulnerability
We analyse how the regression model trained on the ViT-B (Head) data generalizes to other target models. The main points are:
@@
  • From-Scratch-Training: Carlini et al. (2022) provide limited results on from-scratch-training. To the best of our knowledge these are the only published LiRA results on image classification models. Figure˜5(b) displays that our prediction model underestimates the vulnerability of the from-scratch trained target models. We have identified two potential explanations for this: 1 In from-scratch-training all weights of the model need to be trained from the sensitive data and thus potentially from-scratch-training could be more vulnerable than fine-tuning. 2 The strongest attack in Carlini et al. (2022) uses data augmentations to improve the performance. We are not using this optimization. Additionally, as noted in Remark˜3, our theoretical analysis suggests that the power-law would not always hold for general from-scratch training.
@@
4.4 Individual MIA vulnerability
In order to assess the per-sample MIA vulnerability, we run the experiment with $257$ models with each of them once acting as a target and as a shadow model otherwise, compute the tpr at fpr for every sample separately. In Figure˜6, we display the individual vulnerability as a function of $S$ (shots) for Patch Camelyon. The plot shows the maximal vulnerability of all samples and different quantiles over six seeds. The solid line is the median and the errorbars display the min and max over seeds. The quantiles are more robust to extreme outliers and show decreasing trends already at much lower $S$ than the maximum vulnerability.
Figure 6: Individual vulnerability for ViT-B (Head) when fine-tuning on Patch Camelyon. We observe a similar power-law relationship for individuals when looking at the quantiles but the max decreases slower.
@@ 4.5 Comparison between empirical models and universal DP bounds
Using the translation between the $(\textsc{tpr},\textsc{fpr})$ to $(\epsilon,\delta)$-DP guarantees proposed by Kairouz et al. (2015), we compute the minimum $S$ predicted by our empirical models such that the predicted tpr matches the theoretical bound at target DP level. (See Appendix˜E for a more detailed description.) Table˜1 shows the resulting lower bounds of $S$ for various DP levels and values of fpr. While our empirical observations do not provide any formal guarantees, the comparison serves as an illustration to better understand the different requirements of the average and individual vulnerability. We can see that both for the average and the worst case, obtaining a low fpr would require a large amount of samples per class in order to match the tpr for meaningfully strong DP bounds.
@@ 5 Discussion
We mostly consider LiRA in our paper which is optimal for our simplified model of membership inference (Section˜3.4) but for the transfer-learning experiments (Section˜4) there might be stronger attacks in the future. Furthermore, our simplified model assumes well-behaved underlying distributions, meaning that the data is normally distributed around the class centres. We leave the analysis of other data distributions (e.g., heavy-tailed distributions) to future work.. Formal bounds on MIA vulnerability would require something like DP. In addition, both our theoretical and empirical analysis focus on deep transfer learning using fine-tuning. Models trained from scratch are likely to be more vulnerable.
@@ A.4 Offline LiRA
$\log(\textsc{tpr}_{\mathrm{offLiRA}}(\bm{x})-\textsc{fpr}_{\mathrm{offLiRA}}(\bm{x}))\\ =-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{offLiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}+\log(1+\xi(S))$ (A44)
@@
$\log(\textsc{tpr}_{\mathrm{offLiRA}}(\bm{x})-\textsc{fpr}_{\mathrm{offLiRA}}(\bm{x}))\approx-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-m_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}.$ (A45)

### v6

@@ Abstract
Membership inference attacks (MIAs) are used to test practical privacy of machine learning models. MIAs complement formal guarantees from differential privacy (DP) under a more realistic adversary model. We analyze MIA vulnerability of fine-tuned neural networks both empirically and theoretically, the latter using a simplified model of fine-tuning. We show that the vulnerability of non-DP models when measured as the attacker advantage at a fixed false positive rate reduces according to a simple power law as the number of examples per class increases. A similar power-law applies even for the most vulnerable points, but the dataset size needed for adequate protection of the most vulnerable points is very large.
@@ 1 Introduction
To achieve our goal, we theoretically analyze and systematically apply two state-of-the-art black-box MIAs, LiRA (Carlini et al., 2022) and RMIA (Zarifzadeh et al., 2024), to help understand practical privacy risks when fine-tuning deep-learning-based classifiers without DP protections. For the theoretical model that we analyze, LiRA is the optimal attack by the Neyman–Pearson lemma (Neyman and Pearson, 1933). Under the black-box threat model, in which the adversary does not have access to the model parameters, LiRA and RMIA have been shown to empirically outperform other attacks, especially when the number of shadow models is sufficiently large.
@@
Related work There has been evidence that classification models with more classes are more vulnerable to MIA (Shokri et al., 2017), models trained on fewer samples can be more vulnerable (Chen et al., 2020; Németh et al., 2025), and classes with less examples tend to be more vulnerable (Chang and Shokri, 2021; Kulynych et al., 2022; Tonni et al., 2020). A larger generalization error, which is related to dataset size, has also been shown to be sufficient for MIA success (Song and Mittal, 2021), though not necessary (Yeom et al., 2018). Similarly, minority subgroups tend to be more affected by DP (Suriyakumar et al., 2021; Bagdasaryan et al., 2019). Feldman and Zhang (2020) showed that neural networks trained from scratch are required to memorize a significant fraction of their training data to obtain high utility, while the memorization is greatly reduced for fine-tuning. Additionally, Tobaben et al. (2023) reported how the MIA vulnerability of few-shot image classification is affected by the number of shots. Yu et al. (2023) studied the relationship between the MIA vulnerability and individual privacy parameters for different classes. Recently, worst-case MIA vulnerability has gained more attention (Guépin et al., 2024; Meeus et al., 2024; Azize and Basu, 2025). Nonetheless, the prior works do not consider the rate of change in the vulnerability evaluated at a low fpr, as dataset properties change. Our work significantly expands on these works by a) explicitly identifying a quantitative relationship between dataset properties and MIA vulnerability, i.e., the power-law in Equation˜1, and b) focusing on the worst-case vulnerability, both evaluated at a low FPR. This in turn allows us to extrapolate MIA vulnerability to DP guarantee.
@@
    MIA experiments on the average case vulnerability: We conduct a comprehensive study of MIA vulnerability (tpr at a fixed low fpr) in the transfer learning setting for image classification tasks with target models trained using many different datasets with varying sizes and confirm the theoretical power-law between the number of examples per class and the vulnerability to MIA (see Figures˜1 and 4.2). We fit a regression model which follows the functional form of the theoretically derived power-law. We show both a very good fit on the training data as well as a good prediction quality on unseen data from a different feature extractor and when fine-tuning other parameterizations (see Section˜4.3).
@@ 3.4 A simplified model of the optimal membership inference
We sample vectors on a high-dimensional unit sphere and classify them based on inner product with estimated class mean. This model is easier to analyze theoretically than real-world deep learning examples. We generate the data and form the classifiers (which are our target models) as follows:
@@
    For each class, we first sample a true class mean $\bm{m}_{c}$ on a high dimensional unit sphere that is orthogonal to all other true class means ($\forall i,j\in\{1,\dots,C\},i\neq j:\bm{m}_{i}\perp\bm{m}_{j}$).
@@
    We sample $2S$ vectors $\bm{x}_{c}$ for each class. We assume that they are Gaussian distributed around the the true class mean $\bm{x}_{c}\sim\mathcal{N}(\bm{m}_{c},\Sigma)$ where the $\Sigma$ is the in-class covariance.
@@
    For each “target model” we randomly choose a subset of size $CS$ from all generated vectors and compute per-class means $\hat{\bm{m}}_{c}$.
@@
    The computed mean is used to classify sample $\bm{x}$ by computing the inner product $\langle\bm{x},\hat{\bm{m}}_{c}\rangle$ as a metric of similarity.
@@
$\log(\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))\\ =-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-\bm{m}_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}+\log(1+\xi(S)),$ (9)
where $\bm{m}_{\bm{x}}$ is the true mean of class $y$ and $\xi(S)=O(1/\sqrt{S})$. For large $S$ we have
$\log(\textsc{tpr}_{\mathrm{LiRA}}(\bm{x})-\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))\approx-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-\bm{m}_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}.$ (10)
@@
This implies that if $\bm{x}-\bm{m}_{\bm{x}}||$ is bounded, then the worst-case vulnerability is also bounded. Hence we can significantly reduce the MIA vulnerability of all examples in this non-DP setting by simply increasing the number of examples per class.
@@
By Lemma˜1 and the proof of Theorem˜2, a necessary condition for the power-law is that $(\mu_{\mathrm{in}}(\bf{x})-\mu_{\mathrm{out}}(\bf{x}))/\sigma(\bf{x})$ converges to zero at rate $O(1/S^{\alpha})$ with $\alpha>0$. In our simplified model, this holds with $\alpha=1/2$. However, $\mu_{\mathrm{in}}(\bf{x})-\mu_{\mathrm{out}}(\bf{x})\rightarrow 0$ might not always be the case for larger neural networks trained from scratch. Furthermore, even if $(\mu_{\mathrm{in}}(\bf{x})-\mu_{\mathrm{out}}(\bf{x}))/\sigma(\bf{x})$ converges to zero, it is not clear what the convergence rate would be.
@@
$\log(\overline{\textsc{tpr}}_{\mathrm{LiRA}}-\overline{\textsc{fpr}}_{\mathrm{LiRA}})\approx-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\overline{\textsc{fpr}}_{\mathrm{LiRA}})^{2}+\log\left(\mathbb{E}_{(\bm{x},y)\sim{\mathbb{D}}}\left[\frac{|\langle\bm{x},\bm{x}-\bm{m}_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}\right]\right).$ (12)
@@ 4.1 Experimental setup
We report results for fine-tuning a last layer classifier (Head) trained on top of a Vision Transformer ViT-Base-16 (ViT-B; Dosovitskiy et al., 2021), pre-trained on ImageNet-21k (Russakovsky et al., 2015). The results for using ResNet-50 (R-50; Kolesnikov et al., 2020) as a backbone can be found in Section˜D.1. We optimize the hyperparameters (batch size, learning rate and number of epochs) using the Optuna library (Akiba et al., 2019) with the Tree-structured Parzen Estimator (TPE; Bergstra et al., 2011) sampler with 20 iterations (more details in Section˜C.2). We provide the the code for reproducing the experiments in an open repository [^1] https://github.com/DPBayes/impact-dataset-properties-MI-vulnerability-deep-TL.
@@ 4.3 Model to predict dataset vulnerability
We analyze how the regression model trained on the ViT-B (Head) data generalizes to other target models. The main points are:
@@
  • From-Scratch-Training: Carlini et al. (2022) provide limited results on from-scratch-training. To the best of our knowledge these are the only published LiRA results on image classification models. Figure˜5(b) displays that our prediction model underestimates the vulnerability of the from-scratch trained target models. We have identified two potential explanations for this: 1 In from-scratch-training all weights of the model need to be trained from the sensitive data and thus potentially from-scratch-training could be more vulnerable than fine-tuning. 2 The strongest attack in Carlini et al. (2022) uses data augmentations to improve the performance. We are not using this optimization. Additionally, as noted in Remark˜3, our theoretical analysis is based on a simple model and the power-law might not occur at all for larger neural networks trained from scratch.
@@
Figure 6: Individual vulnerability for ViT-B (Head) when fine-tuning on Patch Camelyon. We observe a similar power-law relationship for individuals when looking at the quantiles but the max decreases slower. The Average Case shows the average LiRA vulnerability over the individuals also illustrated in the right panel of Figure˜3.
@@
Figure 7: Illustration of the extrapolation for the average-case at $\textsc{fpr}=10^{-3}$ (left) and worst-case at $\textsc{fpr}=10^{-1}$ (right).
4.4 Individual MIA vulnerability
In order to assess the per-sample MIA vulnerability, we run the experiment with $257$ models with each of them once acting as a target and as a shadow model otherwise, compute the tpr at fpr for every sample separately. In Figure˜6, we display the individual vulnerability as a function of $S$ (shots) for Patch Camelyon and compare it with corresponding average case vulnerability from Figure˜3. The plot shows the maximal vulnerability of all samples and different quantiles over six seeds. The solid line is the median and the errorbars display the min and max over seeds. The quantiles are more robust to extreme outliers and show decreasing trends already at much lower $S$ than the maximum vulnerability.
@@ 4.5 Comparison between empirical models and universal DP bounds
Using the translation between the $(\textsc{tpr},\textsc{fpr})$ to $(\epsilon,\delta)$-DP guarantees proposed by Kairouz et al. (2015), we compute the minimum $S$ predicted by our empirical models such that the predicted tpr matches the theoretical bound at target DP level. We illustrate this in Figure˜7. (See Appendix˜E for a more detailed description.) Table˜1 shows the resulting lower bounds of $S$ for various DP levels and values of fpr. While our empirical observations do not provide any formal guarantees, the comparison serves as an illustration to better understand the different requirements of the average and individual vulnerability. We can see that both for the average and the worst case, obtaining a low fpr would require a large amount of samples per class in order to match the tpr for meaningfully strong DP bounds.
@@ 5 Discussion
We mostly consider LiRA in our paper which is optimal for our simplified model of membership inference (Section˜3.4), but for the transfer-learning experiments (Section˜4) there might be stronger attacks in the future. Furthermore, our simplified model assumes well-behaved underlying distributions, meaning that the data is normally distributed around the class centers. We leave the analysis of other data distributions (e.g., heavy-tailed distributions) to future work. Formal bounds on MIA vulnerability would require something like DP. In addition, both our theoretical and empirical analysis focus on deep transfer learning using fine-tuning. Models trained from scratch are likely to be more vulnerable. Concurrently, a related paper (Hayes et al., 2025) has studied the relation between LiRA vulnerability and dataset size when attacking LLMs during the pre-training phase. Extending this analysis is promising but computationally very expensive.
@@ A.4 Offline LiRA
$\log(\textsc{tpr}_{\mathrm{offLiRA}}(\bm{x})-\textsc{fpr}_{\mathrm{offLiRA}}(\bm{x}))\\ =-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{offLiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-\bm{m}_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}+\log(1+\xi(S))$ (A44)
@@
$\log(\textsc{tpr}_{\mathrm{offLiRA}}(\bm{x})-\textsc{fpr}_{\mathrm{offLiRA}}(\bm{x}))\approx-\frac{1}{2}\log S-\frac{1}{2}\Phi^{-1}(\textsc{fpr}_{\mathrm{LiRA}}(\bm{x}))^{2}+\log\frac{|\langle\bm{x},\bm{x}-\bm{m}_{\bm{x}}\rangle|}{\sqrt{\bm{x}^{T}\Sigma\bm{x}}\sqrt{2\pi}}.$ (A45)